OpenStudy (anonymous):

if x^2+bx=(x+p)^2+q, express p and q in terms of b????

OpenStudy (phi):

The general idea is, if you have \[ x^2 +bx = x^2+2x+c \] you can match coefficients: the b in front of the x on the left side matches the 2 in front of the x on the right side c on the right must match (the unwritten) zero on the left. so you can say b=2, c=0 to do your problem, expand the right side, and match coefficients.

OpenStudy (anonymous):

how does the bx equate to a 2x?

OpenStudy (phi):

I posted an example. It is not your problem. But if we told \[ x^2 +bx = x^2+2x+c \] then we can match coefficients.

OpenStudy (phi):

Have you multiplied out (x+p)^2 out ?

OpenStudy (anonymous):

x^2+2px+p^2.

OpenStudy (phi):

so what is your problem now? x^2+bx= x^2 +2px +p^2+q can you match coefficients?

OpenStudy (anonymous):

my problem is still the same.

OpenStudy (phi):

Maybe this idea does not make sense, but what you do for this problem is match the coefficient of the x term on the left with the coefficient of the x term on the right.

OpenStudy (anonymous):

what form is the first equation written in?

OpenStudy (phi):

x^2+bx= x^2 +2px +p^2+q

OpenStudy (anonymous):

no the stuff on the LHS.

OpenStudy (phi):

I am not following your question. But for this problem: if x^2+bx=(x+p)^2+q, express p and q in terms of b???? we expand the (x+p)^2 so that we get x^2 and x's on the right side: x^2+bx= x^2 +2px +p^2+q now we match coefficients of the x term and the constant term

OpenStudy (anonymous):

so the x^2s cancel out

OpenStudy (phi):

We are not "solving" in the normal sense. We are doing something slightly easier: the numbers in front of the x^2 term, the x term, and the x^0 (constant term) must match to make the left equal to the right side.

OpenStudy (phi):

so 1*x^2 on both sides is ok. but to make the left side equal the right, the b must match the 2p.

OpenStudy (anonymous):

ok sorry I do not get this

OpenStudy (anonymous):

thanks for trying to help me.

OpenStudy (phi):

Think about this example 3x= bx for all x. what is b? Not 1: 3x≠1x, (this is only true for x=0, but we want it true for any x). Think about it, and you see b=3 makes both sides equal for any x.

OpenStudy (anonymous):

ok then

OpenStudy (phi):

The is the first idea. The second idea is 2x^2 = bx what b makes both sides equal (for all x)? b is a number, and no matter what number we can never get this to work for all x. for example 2x^2≠ 2x if x=3 (2*9≠2*3) the only way to get it to work is if we started with 2x^2= bx^2 (x^2 on both sides)

OpenStudy (phi):

so if we started with 2x^2 + 3x= ax^2 +bx the only way to make both sides the same is have a=2 and b=3 we match coefficients

OpenStudy (anonymous):

it makes a little sense.

OpenStudy (phi):

First, finish off the problem, then think about it later (sometimes these things make sense after a rest) x^2+bx= x^2 +2px +p^2+q b= 2p an p^2+q= 0 we can think of the left side as x^2+bx+0*x^0 and the right side as x^2 +2px +(p^2+q)x^0 (of course x^0 is 1)

OpenStudy (phi):

p in terms of b, we get from b=2p solve for p: p= b/2 now use p^2+q=0 q= -p^2 replace p with b/2 to get q in terms of b

OpenStudy (phi):

Maybe your intuition will work if you think of each power of x (x^2 , x^1, x^0) as a different animal, and you are told alice and bob have the same number and types of animals: alice has 2 cows and 3 horses. bob has a cows and b horses. what is a and b? now replace cows with x^2 and horses with x^1 same idea